# Learning Koopman Dynamics for Safe Legged Locomotion with Reinforcement Learning-based Controller

### v1

@@ VI Conclusion and Future Work
In this work, we present a novel safe navigation framework for legged robots by combining the Koopman operator theory and the MPC framework. We first effectively learns the dynamics from the collected trajectories using Dynamic Mode Decomposition without the burden of hyperparameter tuning. Then, we develop an MPC framework for safe navigation that efficiently solves the given constrained optimization problem by leveraging the linearity of the learned Koopman dynamics. We show that the Koopman dynamics, especially with Time-delay Embedding, offers improved accuracy for the long-term forward dynamics prediction task. We also deploy our safe navigation framework with the Unified Linear model in environments with narrow passages and mazes and demonstrate improved safety compared to the baselines.
There are several directions for future work. For instance, the utility of the learned linear dynamics is not limited to obstacle avoidance discussed in this work. There exists a wide range of potential tasks, including whole-body control of legged robots, loco-manipulation, or traversing more complex environments. These tasks require modeling the full-body dynamics, which likely requires a careful investigation of the lifting function for the reliable Koopman linearization. Once the Koopman dynamics is learned, it will be possible to leverage similar model-based safety frameworks. We also aim to extend the capability of this method by exploring diverse lifting techniques. One possibility is to utilize neural network-based lifting functions, which could potentially provide a more concise and accurate dynamics model. However, this may introduce complexity to the implementation of the MPC framework. Investigating this trade-off will be an interesting future research direction for legged robots.

### v2

@@ VI Conclusion and Future Work
In this work, we present a novel safe navigation framework for bipedal robots by combining Koopman operator theory with the MPC framework. We first train a low-level locomotion control policy using deep RL. Then, we learn the nonlinear base-level dynamics from the collected trajectories using Dynamic Mode Decomposition. Finally, we develop an MPC framework for safe navigation that efficiently solves the given constrained optimization problem by leveraging the linearity of the learned Koopman dynamics. We show that the learned Koopman dynamics offer improved accuracy for long-term forward dynamics prediction. We also deploy our safe navigation framework in challenging environments with narrow passages and mazes and demonstrate improved safety compared to the baselines.
There are several directions for future work. For instance, the utility of the learned linear dynamics is not limited to obstacle avoidance discussed in this work. There exists a wide range of potential tasks, such as loco-manipulation, or traversing more complex environments, which require modeling full-body dynamics. This goal will likely require a careful investigation of the lifting function for the reliable Koopman linearization. We also aim to extend the capability of this method by exploring diverse lifting techniques. One possibility is to utilize time-delay or recurrent neural network based lifting, which could provide a more concise and accurate forward dynamics model. However, this may introduce additional complexity into the implementation of the learning framework. Investigating this trade-off will be an interesting future research direction for legged robots. Finally, the learned linearized model may show severe performance degradation in unseen environments and lead to the failure of safe navigation. It will be interesting to combine our framework with additional safety mechanisms, such as control barrier functions under state uncertainty [47] or additional recovery controllers [48].